OpenStudy (anonymous):

Please help me solve the algebra word problem: The Wilson family and the Alexander family each used their sprinklers last summer. The water output rate for the Wilson family’s sprinkler was 30L per hour. The water out put rate for the Alexander family’s sprinkler was 25L per hour. The families used their sprinklers for a combined total of 45 hour, resulting in a total output of 1200L. How long was each sprinkler used? Wilson family sprinkler = _______ hours Alexander family’s sprinkler = ______ hours

OpenStudy (anonymous):

If you let w = the number of hours the Wilson sprinkler was used, then 45 - w = the number of hours the Alexander sprinkler was used. 30w + 25(45 - w) = 1200 So, now, just solve for "w". Are you able to do that?

OpenStudy (anonymous):

No, I tried and my answers were wrong, I do not understand the substitution theory.

OpenStudy (anonymous):

30w + 25(45 - w) = 1200 -> 30w - 25w = 1200 - (25)(45) (30 - 25)w = 75 -> 5w = 75 -> w = 15 So, the Wilson's usage was 15 hours while the Alexander's usage was 30. In that initial equation: 30w + 25(45 - w) = 1200 this is actually: [30 (lit/hr)](w hr) + [25 (lit/hr)][(45 - w)hr] = 1200 lit -> "hrs" cancels And the Alexander's usage is 45 - w because both usages have to add to 45

OpenStudy (anonymous):

What you have is actually 2 equations in 2 unknowns: 30w + 25a = 1200 and w + a = 45 a = 45 - w (from the second equation) so you take that "45 - w" and put THAT in the first equation where you see the "a".

OpenStudy (anonymous):

The example shows what your equation shows, and I understand what is written but not how you reached the answers.how did you get 75? Just viewed your last statement I understand what you have written but not how your received the answer.

OpenStudy (anonymous):

I am asking exactly how to reach the solution, is it by dividing a then multiplying it by w?

OpenStudy (anonymous):

The Wilson family used more water why is the number of hours less than the Alexander family?

OpenStudy (anonymous):

The 75 came from: 1200 - (25)(45) -> 1200 - 1125 -> 75 Here, let me fill in a few more steps: 30w + 25(45 - w) = 1200 30w + (25)(45) - 25w = 1200 30w - 25w + (25)(45) = 1200 (30 - 25)w = 1200 - (25)(45) 5w = 1200 - 1125 = 75 5w = 75 -> w = 75 / 5 = 15

OpenStudy (anonymous):

I understand that

OpenStudy (anonymous):

So, you're all set?

OpenStudy (anonymous):

Wouldn't the 15 belong to the Alexander family since we received that answer from using the 25 L?

OpenStudy (anonymous):

No, because in my first post, we established that "w" (w for Wilson, but we could have used any variable) is the # of hours of Wilson's usage. So, when we solved for "w", that value pertains to the Wilsons.

OpenStudy (anonymous):

At this point, you might want to step away from "how to solve it" and just plug the values back in to see how it does indeed work.

OpenStudy (anonymous):

T.C. Thank you so very much. You deserve a medal, please tell me how to give you a medal, this is just my second time on this site.

OpenStudy (anonymous):

oh, ok, you just hit "best response" (the blue box) by my name. And it was very nice working with you!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Tuiti You are very welcome for the help!

OpenStudy (anonymous):

Just gave you a medal and it is so nice to have someone explain what seems to be Greek to me. Thank you again!

OpenStudy (anonymous):

Before you go, this was a good problem for "2 equations in 2 variables". We went through it in fairly big detail, so you or anyone would benefit from just going over it at least one more time.

OpenStudy (anonymous):

I hope to work with you again, thanks for being here for us!

OpenStudy (anonymous):

You made my day! @Tuiti

OpenStudy (anonymous):

LOL! You made mine too!

OpenStudy (anonymous):

T.C can you help me to solve the compound inequality. Both signs are less than or equal to, it sometimes does not show up when I copy and paste. 4y- 2 ≤ 22 or 3y-5 ≤ 1 Thanks Tuiti

OpenStudy (anonymous):

T.C can you help me to solve the compound inequality. Both signs are less than or equal to, it sometimes does not show up when I copy and paste. 4y- 2 ≤ 22 or 3y-5 ≤ 1 The answer has to be in interval notation.